How can I go prove ( sqrt( (n+1)^3 ) ) is Big Omega (n * sqrt(n)) using the formal definitions of Big Oh, Big Theta, and Big Omega?
Answer #1
f(n) = Omega(g(n)) means there are positive constants c and n0, such that f(n) >= cg(n) for all n ≥ n0

sqrt((n+1)^3) => (n+1) sqrt(n+1)
(n+1) sqrt(n+1) = Omega(n * sqrt(n))

=>  (n+1) sqrt(n+1) >= cn * sqrt(n)
Let's assume c = 1
=>  (n+1) sqrt(n+1) >= n * sqrt(n)
=>  (n+1) >= n  (because sqrt(n+1) is always > sqrt(n))  
(n+1) >= n for all positive values of n

so, (n+1) sqrt(n+1) = Omega(n * sqrt(n)) for c = 1 and n0 = 0